36 ml of pure water takes 100 second to evaporate from a vessel and ETA connected to an electric source of which delivers 806 watt the delta H of H2O is

Grade:12

1 Answers

Satvik
5 years ago
Volume of water= 36mlWe know that Density= M/V1= M/36 ( Density of water = 1g/ml )M=36 gm Moles= mass/ molar mass of water n=2 molNow Energy= Power × time=806 × 100= 80600 J =80.6kJ2mol= 80.6kJ 1mol=80.6/2 = 40.3 kJ (Ans.)